OpenStudy (anonymous):

Can someone please show me how they do this? 3(x – 2) + 4x = -27

OpenStudy (anonymous):

3(x-2) + 4x= -27 3x-6+4x=-27 7x-6=-27 +6 +6 7x=-27 then divide both sides by 7 x=?

OpenStudy (anonymous):

x=-3.85?

OpenStudy (anonymous):

oh sorry 7x=-21 then solve for x x=-3

OpenStudy (anonymous):

You're brilliant! Thank you girl :)

OpenStudy (anonymous):

No problem you can ask me more questions if you want. i have one open that i don't know... maybe you could help me out too